In 2019, X Company's profit function was 0.31R - $90,500, where R is revenue. In 2020, the relationship between revenue and variable costs will not change, but fixed costs will increase by $17,195. Assuming a tax rate of 40%, what will revenue have to be in order for X Company to earn $46,400 after taxes in 2020?
Answer #1

Answers :$ 596.865.59 (or) 596,864.52

Contibution Margin Ratio = 31%

Fixed cost = 90,500+17195 = 107,695

Net income Required = 46,400

Net operating income Must ear = 46,400 /(1-TAx) = 46,400 /60% = 77,333.33

REvenue required = (Fixed cost + Revenue Required ) / CM Ratio

= (107,695+77,333.00 )/31%= 185,028 /31% = $ 596.865.59 (or) 596,864.52